reasoning we are two numbers in a fact family we have a difference of 1 and a sum of 13. What numbers are we? how do you know?
Answer by ankor@dixie-net.com(22740)

we have a difference of 1 and a sum of 13.
a - b = 1
a + b = 13
------------adding eliminates b, find a
2a = 14
a = 7
and
7 + b = 13
b = 13 - 7
b = 6
:
What numbers are we? 7 & 6
:
Check values in both equations, then you know